OpenStudy (anonymous):

Calcium carbide (CaC2) is manufactured by combining lime (CaO) and carbon at high temperatures, as shown in the equation. Approximately how many grams of CaC2 can be formed from 84.0 g of CaO and 48.0 g of C?CaO(s) + 3C(s) CaC2 + CO(g) A.132 g B.96.0 g C.64.0 g D.85.4 g

OpenStudy (anonymous):

Do you know how to find the limiting reactant?

OpenStudy (anonymous):

@LeibyStrauss I have no idea how to go about this problem. I chose B but it was wrong. So I want to learn how to do this right.

OpenStudy (anonymous):

Suppose I eat 1 trail mix (a day) of 2 almonds and 7 walnuts. If I have 100 almonds 70 walnuts, which nut will I have less of? Another way of asking this is, which nut will limit the number of trail mixes that I can make?

OpenStudy (anonymous):

the 70 walnuts? @LeibyStrauss

OpenStudy (anonymous):

Yes. Mathematically we can solve it as follows: For every 2 almonds we need 7 walnuts So we can write either ratio below: \[\frac{ 2 almonds }{ 7 walnuts} or \frac{ 7 walnuts }{ 2 almonds } \] So if we want to know how many walnuts is needed if we have 100 almonds we can write the equation below. Since we are writing almonds first (which is in the numerator) we will write the next ratio with almonds in the denominator, so that almonds and almonds cancel and we are left with walnuts. \[\frac{ 100 almonds}{1 }*\frac{ 7 walnuts}{ 2almonds } = 350 walnuts\] This means that 350 walnuts are needed. Since we only have 70 walnuts, walnuts is the limiting ingredient. If this makes sense we can use this method for you question.

OpenStudy (anonymous):

Thankyou soooooo much :) @LeibyStrauss Your explanation was awesome

OpenStudy (anonymous):

Did you find the limiting reactant in your problem?